OpenStudy (anonymous):

Write an equation of a line that passes through the point (7, 3) and is parallel to the line y = -2/3 x + 3 A. y = 3/2 x -3 B. y = 3/2 x +3 C. y = -2/3 x + 23/3 D. y= -2/3 x - 23/3

OpenStudy (anonymous):

@Abmon98

OpenStudy (texaschic101):

do you know what the slope is in the equation y = -2/3x + 3 ?

OpenStudy (anonymous):

-2/3 correct? @texaschic101

OpenStudy (texaschic101):

correct. A parallel line will have the same slope. now use y = mx + b slope(m) = -2/3 (7,3) x = 7 and y = 3 now we sub and solve for b 3 = -2/3(7) + b 3 = -14/3 + b 3 + 14/3 = b 9/3 + 14/3 = b 23/3 = b so your parallel equation is : y = -2/3x + 23/3 any questions ?

OpenStudy (anonymous):

Oh yeah! That makes sense. I forgot that all parallel lines have the same slope. That's really easy, basically just plugging in numbers. My brain really doesn't work in the mornings -_- Thank you! @texaschic101

OpenStudy (texaschic101):

lol...my brain sometimes does and sometimes doesn't. This morning, my brain seems to be doing ok :)